xx Obtaining Reports of Timer Information

Obtaining Reports of Timer Information

Cactus Maintainers <cactusmaint@cactuscode.org>

July 06 2003

Abstract

This thorn provides mechanisms for obtaining different information from timers during a Cactus run. Currently there are only simple reports to screen but in the future more configurable options will be added.

1 Using This Thorn

To use this thorn you only need to include it in your ActiveThorns parameter and set the relevent parameters. With default parameters the thorn does nothing.

Current options are to output the timing report to screen at the next iteration (via a steering interface), before each checkpoint, at a given iteration, or at periodic iteration intervals.

1.1 Obtaining This Thorn Available as part of the Cactus Computational Toolkit in arrangement CactusUtils.

1.2 Support and Feedback To the Cactus Team at cactusmaint@cactuscode.org or through the Bug/Feature Request web pages at

www.cactuscode.org.

2 Parameters

11 allxtimersxclock 10 Scope: private01 STRING

11 Description: Which clock to use for the all timers output

11 Range 10 01 Default: gettimeofday
11 .*  10 any clock name allowed

11

11 beforexcheckpoint  10 Scope: private01  BOOLEAN

11 Description: Before a checkpoint

11                       10              01  Default: no

11

11 nxtopxtimers  10 Scope: private01  INT

11 Description: How many timers to include in the top timer report

11 Range 10 01 Default: (none)
11                       10 Do not print the report
11 1:*  10 Any number of timers

11

11 next  10 Scope: private01  BOOLEAN

11 Description: On next iteration

11                       10              01  Default: no

11

11 outxat  10 Scope: private01  INT

11 Description: Output timer information at a given iteration

11 Range 10 01 Default: -1
11 -1  10 Do not output at specific iteration (default)
11 0:*  10 At this iteration

11 outxevery  10 Scope: private01  INT

11 Description: How often to output timer report to screen

11 Range 10 01 Default: (none)
11                       10 No periodic output (default)
11 1:*  10 Every so many iterations

11

11 outxfilename  10 Scope: private01  STRING

11 Description: File name for timer reports

11 Range 10 01 Default: (none)
11 ˆ$  10 empty filename: print to stdout
11 ˆ. +$  10 otherwise: print to that file

11

11 outputxallxtimers  10 Scope: private01  BOOLEAN

11 Description: Output one file per processor containing all the Cactus timers

11                       10              01  Default: no

11

11 outputxallxtimersxreadable 10 Scope: private01  BOOLEAN

11 Description: Output one file containing information about all the Cactus timers (average, minimum, and maximum over all processors), in a format that is readable by humans

11                       10              01  Default: no

11

11 outputxallxtimersxtogether 10 Scope: private01  BOOLEAN

11 Description: Output three files (formats .txt, .csv, and .tsv), containing information about all the Cactus timers (average, minimum, and maximum over all processors)

11                       10              01  Default: no

11

11 outputxschedulextimers 10 Scope: private01  BOOLEAN

11 Description: Output the schedule timers in a formatted tabular format

11                       10              01  Default: yes

11

11 checkpointxevery  10 Scope: shared from IO01INT

11

11 outxdir  10 Scope: shared from IO01STRING

3 Interfaces

General Implements: timerreport

4 Schedule This section lists all the variables which are assigned storage by thorn CactusUtils/TimerReport. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxANALYSIS x timerreportxoutputevery x print the timer report x

11 x 11 Language: 11 c
11 x 11 Options: 11 global
11 x 11 Type:  11 function

11

CCTKxTERMINATE x timerreportxoutputterminate x print the timer report x

11 x 11 Before: 11 driverxterminate
11 x 11 Language: 11 c
11 x 11 Options: 11 global
11 x 11 Type:  11 function

CCTKxCHECKPOINT x timerreportxcheckpoint x print the timer report x

11 x 11 Language: 11 c
11 x 11 Options: 11 global
11 x 11 Type:  11 function

11

Aliased Functions x

11 Alias Name: xxxxxxx  11 Function Name:
11 TimerReportxCheckpoint  11 zzzxTimerReportxCheckpoint
11 TimerReportxOutputEvery  11 zzzxTimerReportxOutput
11 TimerReportxOutputTerminate 11 zzzxTimerReportxOutput